# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ 6.5 Brain-inspired OWL
Compared to DNN based artificial intelligence systems, biological organisms like humans naturally have the strong ability of open-world learning, enabling them to interact with the environment throughout their lifetime [160]. For example, humans can be aware of when they are likely to be wrong and then refuse to make a high-risk decision; they can discover novel knowledge based on what they have learned previously, and last but not least, humans excel at learning from a dynamically changing environment incrementally without suffering from catastrophic forgetting In light of this, some works [123, 124, 204, 205, 206] aim to design brain-inspired open-world machine learning models. For example, van de Ven et al. [123] proposed a brain-inspired feature replay method that uses the original classification model itself to generate features of old classes, avoiding using additional generative models and showing promising performance for CIL. Recently, inspired by biological learning systems, Wang et al. [204] built a continual model with the ability of stability protection and active forgetting. We believe that the biological cognitive system will inspire more advanced OWL models to narrow the gap between artificial and natural intelligence.
6.6 Open-world Unlearning
@@
Most of the focus in machine learning is on training models to acquire new knowledge, e.g., incremental learning aims to incorporate new knowledge without forgetting previously learned ones. On the contrary, there are scenarios where machine unlearning may be required for various reasons: (1) Sensitive information removal. In some cases, machine learning models may inadvertently memorize sensitive or private information from the training data [207]. Unlearning can be used to erase such information from the model’s memory to ensure data privacy and compliance with regulations. (2) Bias mitigation. If a model exhibits undesirable biases in its predictions [208], unlearning can be employed to eliminate these biases or align the model better with desired behaviors by removing some patterns learned during training. (3) Selective forgetting. Unlearning allows models to discard some outdated information to leave more room for learning new tasks. This is different from the well-known catastrophic forgetting [121, 14] which may forget important information that we want to memorize. Indeed, machine unlearning is challenging because it involves determining which parts of knowledge to remove, developing algorithms for knowledge erasure, and ensuring that unlearning does not adversely affect the model’s overall performance [209, 210]. Therefore, it is a long-term research goal to perform open-world learning and unlearning flexibly and simultaneously.
7 Conclusion
@@
Open-world learning stands as a critical frontier in the field of artificial intelligence, particularly as AI systems increasingly operate in complex, real-life environments (e.g. self-driving cars, mobile robots, chatbots, etc.), and engage with humans and automated systems. The essence of open-world learning lies in the ability to detect and adapt to new diverse and ever-changing scenarios, like a self-motivated learner. In this article, we provided an overview of open-world learning, with an emphasis on techniques concerning unknown rejection, novel class discovery, and class-incremental learning. Finally, we outlined several challenging open problems that merit substantial research endeavors in the future.

### v3

@@ VI Future Directions
5) Brain-inspired Open-world Learning. Compared to DNN-based artificial intelligence systems, biological organisms like humans naturally have the strong ability of open-world learning, enabling them to interact with the environment throughout their lifetime [361]. For example, humans can be aware of when they are likely to be wrong and then refuse to make a high-risk decision; they can discover novel knowledge based on what they have learned previously, and last but not least, humans excel at learning from a dynamically changing environment incrementally without suffering from catastrophic forgetting In light of this, some works [159, 119, 125, 424, 425] aim to design brain-inspired open-world machine learning models. For example, van de Ven et al. [159] proposed a brain-inspired feature replay method that uses the original classification model itself to generate features of old classes, avoiding using additional generative models and showing promising performance for CL. Recently, inspired by biological learning systems, Wang et al. [125] built a continual model with the ability of stability protection and active forgetting. Inspired by human memory, Liu et al. [426] introduced a biomimetic continual learning framework that combines semi-parametric memory with a wake–sleep consolidation mechanism, enabling deep neural networks to acquire new tasks without forgetting prior knowledge. Building on these advances, future research in brain-inspired open-world learning should focus on developing mechanisms for adaptive confidence estimation to gauge when models should seek feedback [427], fostering active exploration and curiosity-driven modules to autonomously discover novel information [428], integrating multimodal data streams to form richer knowledge representations. Besides, drawing inspiration from the brain’s sparse, it is important to create energy-efficient edge implementations and adaptation in ever-changing environments, ultimately narrowing the gap between artificial and natural intelligence.
@@
6) Open-world Unlearning. Most of the focus in machine learning is on training models to acquire new knowledge, e.g., continual learning aims to incorporate new knowledge without forgetting previously learned knowledge. On the contrary, there are scenarios where machine unlearning may be required for various reasons: (1) Sensitive information removal. In some cases, machine learning models may inadvertently memorize sensitive or private information from the training data [429, 430]. Unlearning can be used to erase such information from the model’s memory to ensure data privacy and compliance with regulations. (2) Bias mitigation. If a model exhibits undesirable biases in its predictions [431], unlearning can be employed to eliminate these biases or align the model better with desired behaviors by removing some patterns learned during training. (3) Selective forgetting. Unlearning allows models to discard some outdated information to leave more room for learning new tasks. This is different from the well-known catastrophic forgetting [340, 235], which may forget important information that we want to memorize. Indeed, machine unlearning is challenging because it involves determining which parts of knowledge to remove, developing algorithms for knowledge erasure, and ensuring that unlearning does not adversely affect the model’s overall performance [432, 433]. Therefore, it is a long-term research goal to perform open-world learning and unlearning flexibly and simultaneously. To reach this goal, future research must address several open problems. First, we need principled frameworks to quantify and localize the knowledge to be unlearned without requiring full retraining, especially under resource constraints. In light of this, Zhao et al. [434, 435] defined the problem of continual forgetting, which focuses on selective and progressive removal of information from pre-trained vision models, and propose a method to effectively forget targeted knowledge with minimal impact on retained knowledge. Second, it is crucial to design unlearning methods that are robust to the uncertainty and dynamics of open-world environments, where new classes may appear and old ones may evolve or disappear. Third, the interplay between learning and unlearning processes should be better understood, e.g., how they influence each other and how to balance them effectively to maintain model stability and adaptability.
VII Conclusion
@@
Open-world learning stands as a critical frontier in the field of artificial intelligence, particularly as AI systems increasingly operate in complex, real-life environments, and engage with humans and automated systems. The essence of open-world learning lies in the ability to detect and adapt to new, diverse and ever-changing scenarios, like a self-motivated learner. The paper presents a comprehensive taxonomy, reviews numerous papers, and provides insights into open-world learning, with an emphasis on techniques concerning unknown rejection, novel class discovery, and continual learning. It also offers performance comparison. Finally, we outlined several challenging yet important research directions that merit substantial research endeavors in the future.